# QueryCAD: Grounded Question Answering for CAD Models

### v3

@@ Abstract
CAD models are widely used in industry and are essential for robotic automation processes. However, these models are rarely considered in novel AI-based approaches, such as the automatic synthesis of robot programs, as there are no readily available methods that would allow CAD models to be incorporated for the analysis, interpretation, or extraction of information. To address these limitations, we propose QueryCAD, the first system designed for CAD question answering, enabling the extraction of precise information from CAD models using natural language queries. QueryCAD incorporates SegCAD, an open-vocabulary instance segmentation model we developed to identify and select specific parts of the CAD model based on part descriptions. We further propose a CAD question answering benchmark to evaluate QueryCAD and establish a foundation for future research. Lastly, we integrate QueryCAD within an automatic robot program synthesis framework, validating its ability to enhance deep-learning solutions for robotics by enabling them to process CAD models. https://claudius-kienle.github.com/querycad
@@ I Introduction
In the industrial sector, many workflows are centered around Computer-Aided Design (CAD) models. These models contain precise representations of individual parts and their assembly into larger components. Engineers in various industrial domains, such as robot programming, rely heavily on CAD models to extract measurements, identify specific features, and understand how parts interact within a larger system. To automate these still largely manual engineering processes, it is crucial to develop methods that can retrieve and interpret information from CAD models in an automated manner, mirroring the way users interact with them. However, since no web-scale dataset for industrial CAD models exists, it is generally hard to train or extend existing models to understand this modality [1]. Moreover, there currently exists no method for automatic information retrieval from CAD models. This gap is particularly evident in the context of automated robot program synthesis for industrial applications, where extracting parameters from CAD models is essential.
@@ III Methods
We propose QueryCAD, a system to answer free-form questions about CAD models and their parts. In doing so, QueryCAD accesses SegCAD, a model for segmenting parts of a CAD model matching an open-vocabulary part description. The architecture can be seen in Figure 2. To evaluate QueryCAD, we develop the first benchmark for the task of CAD question answering.
@@ III-A3 Image to CAD Model Alignment
Figure 3: Pruning CAD faces to retain only neighboring ones. From raw faces (2) derived from a 2D mask (1), ray casting identifies the closest face (3). Adjacent faces are then selected recursively via the adjacency graph (4), discarding non-adjacent ones, i.e. ones visible through holes (1). The final pruned part (5) excludes the faces on the right, as they are distant from the viewport (3) and not adjacent to the masked faces on the left.
@@ III-B QueryCAD: CAD Understanding, Reasoning and Answering
To achieve these properties, we implemented the following approach: First, the user’s query is passed to a code-writing LLM using a carefully crafted prompt [^1] See https://claudius-kienle.github.com/querycad for prompts, CAD-Q&A benchmark, and examples.. The objective of the LLM is to generate Python code that computes the measurements that answer the question. For this, the prompt defines Python classes for the CAD model and the CAD part. The Python classes have attributes such as extents or the center position, which can be accessed in the code generated by the LLM. To enable the LLM to search in the CAD model for a specific feature or part, SegCAD is integrated via a Python Application Programming Interface (API). The API calls are parameterized by the LLM with the free-text part description and potential viewing angles.
@@ IV-A Evaluation
Syntax addresses samples where the LLM generated invalid Python syntax, which was the case for none of the samples. Reasoning corresponds to samples where the proposed Python code features incorrect reasoning or hallucinations, which occurred once. If the Python code is valid, there are still two issues that can cause invalid answers: Masks, where incorrect parts were returned by SegCAD due to imprecise, missing or incorrect masks proposed by GroundedSAM (see Fig. 5 top, left) or due to inaccuracies in the ray casting, which occurred 52 times. Lastly, CAD-Interface classifies answers where the CAD interface provided in the Python code was not powerful enough to answer the question, which was the case for 14 samples.
@@ IV-B Ablations
We evaluated SegCAD with two different viewing angle configurations: 6 viewing angles along the main axes and 8 viewing angles from all 8 corners of the CAD model with an azimuth and elevation of the camera of 45 degrees. Figure 5 compares the two viewing configurations. While the configuration with 6 viewing angles makes it harder for GroundedSAM [36] to understand the depth of the CAD model, as most of the features are orthogonal to the viewport, it effectively allows to detect features like through-holes or other geometrical features specifically from that axis. The configuration with 8 viewing angles has the benefit to capture the 3D shape of the model more accurately, reducing the likelihood of wrong predictions by GroundedSAM. One example are shafts and holes, which look the same from the top. However, we observed that with 8 viewing angles, GroundedSAM occasionally misses certain parts, as shown in Figure 5. Additionally, it becomes more challenging to select the parts visible from a specific viewing direction with 8 viewing angles.

### v4

@@ Abstract
CAD models are widely used in industry and are essential for robotic automation processes. However, these models are rarely considered in novel AI-based approaches, such as the automatic synthesis of robot programs, as there are no readily available methods that would allow CAD models to be incorporated for the analysis, interpretation, or extraction of information. To address these limitations, we propose QueryCAD, the first system designed for CAD question answering, enabling the extraction of precise information from CAD models using natural language queries. QueryCAD incorporates SegCAD, an open-vocabulary instance segmentation model we developed to identify and select specific parts of the CAD model based on part descriptions. We further propose a CAD question answering benchmark to evaluate QueryCAD and establish a foundation for future research. Lastly, we integrate QueryCAD within an automatic robot program synthesis framework, validating its ability to enhance deep-learning solutions for robotics by enabling them to process CAD models. https://claudius-kienle.github.io/querycad
[…]
@@ I Introduction
In the industrial sector, many workflows are centered around Computer-Aided Design (CAD) models. These models contain precise representations of individual parts and their assembly into larger components. Engineers in various industrial domains, such as robot programming, rely heavily on CAD models to extract measurements, identify specific features, and understand how parts interact within a larger system. To automate these still largely manual engineering processes, it is crucial to develop methods that can retrieve and interpret information from CAD models in an automated manner, mirroring the way users interact with them. However, since no web-scale dataset for industrial CAD models exists, it is generally hard to train or extend existing models to understand this modality [1]. Moreover, there currently exists no method for automatic information retrieval from CAD models. This gap is particularly evident in the context of automated robot program synthesis for industrial applications, where extracting parameters from CAD models is essential and done manually.
@@ III Methods
We propose QueryCAD, a system to answer free-form questions about CAD models and their parts. To acomplish this, QueryCAD utilizes SegCAD, a model for segmenting parts of a CAD model matching an open-vocabulary part description. The architecture can be seen in Figure 2. To evaluate QueryCAD, we develop the first benchmark for the task of CAD question answering.
@@ III-A3 Image to CAD Model Alignment
Figure 3: Pruning CAD faces to retain only neighboring ones. From raw faces (2) derived from a 2D mask (1), ray casting identifies the closest face (3). Adjacent faces are then selected recursively via the adjacency graph (4), discarding non-adjacent ones, i.e. visible through holes (1). The final pruned part (5) excludes the faces on the right, as they are distant from the viewport (3) and not adjacent to the masked faces on the left.
@@ III-B QueryCAD: CAD Understanding, Reasoning and Answering
To achieve these properties, we implemented the following approach: First, the user’s query is passed to a code-writing LLM using a carefully crafted prompt [^1] See https://claudius-kienle.github.io/querycad for prompts, CAD-Q&A benchmark, and examples.. The objective of the LLM is to generate Python code that computes the measurements that answer the question. For this, the prompt defines Python classes for the CAD model and the CAD part. The Python classes have attributes such as extents or the center position, which can be accessed in the code generated by the LLM. To enable the LLM to search in the CAD model for a specific feature or part, SegCAD is integrated via a Python Application Programming Interface (API). The API calls are parameterized by the LLM with the free-text part description and potential viewing angles.
@@ IV-A Evaluation
Syntax addresses samples where the LLM generated invalid Python syntax, which was the case for none of the samples. Reasoning corresponds to samples where the proposed Python code features incorrect reasoning or hallucinations, which occurred once. If the Python code is valid, there are still two issues that can cause invalid answers: Masks, where incorrect parts were returned by SegCAD due to imprecise, missing or incorrect masks proposed by GroundedSAM (see Fig. 5 top, left) or due to inaccuracies in the ray casting, which occurred 52 times. Lastly, CAD-Interface classifies answers where the CAD interface provided in the Python code lacks the necessary capabilities to adequately address the question, which was the case for 14 samples.
@@ IV-B Ablations
We evaluated SegCAD with two different viewing angle configurations: 6 viewing angles along the main axes and 8 viewing angles from all 8 corners of the CAD model with an azimuth and elevation of the camera of 45 degrees. Figure 5 compares the two viewing configurations. While the configuration with 6 viewing angles makes it harder for GroundedSAM [36] to understand the depth of the CAD model, as most of the features are orthogonal to the viewport, it effectively allows detecting features like through-holes or other geometrical features specifically from that axis. The configuration with 8 viewing angles has the benefit to capture the 3D shape of the model more accurately, reducing the likelihood of wrong predictions by GroundedSAM. One example are shafts and holes, which look the same from the top. However, we observed that with 8 viewing angles, GroundedSAM occasionally misses certain parts, as shown in Figure 5. Additionally, it becomes more challenging to select the parts visible from a specific viewing direction with 8 viewing angles.